OpenStudy (anonymous):

HURRRY & ANSWER PLZZ ....An airline claims that 90% of the time, its planes are on schedule. If three flights are selected at random, what is the probability that all three are on schedule.

OpenStudy (neer2890):

P(planes are on schedule)=0.9 P(planes aren't on schedule)=1-0.9=0.1 P(all 3 planes are on schedule)=0.9*0.9*0.9=0.729